Find the dimensions of a,b and c if a is distance g is gravity due to acceleration and t is time s=at+bg+ct2

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
s=at+bg+ct2 Dimension of at = dimension of bg dimension of at = m.sSo b(m/s2) = m.sSo dimension of B = s^3Now dimension of at = dimestion of ct^3So m.s= c.s^3So dimension of c is m/s2